### v3

# AIRS-Bench: a Suite of Tasks for Frontier AI Research Science Agents

Abstract
LLM agents hold significant promise for advancing scientific research. To accelerate this progress, we introduce AIRS-Bench (the AI Research Science Benchmark), a suite of 20 tasks sourced from state-of-the-art machine learning papers. These tasks span diverse domains, including language modeling, mathematics, bioinformatics, and time series forecasting. AIRS-Bench tasks assess agentic capabilities over the full research lifecycle—including idea generation, experiment analysis and iterative refinement—without providing baseline code. The AIRS-Bench task format is versatile, enabling easy integration of new tasks and rigorous comparison across different agentic frameworks. We establish baselines using frontier models paired with both sequential and parallel scaffolds. Our results show that agents exceed human SOTA in four tasks but fail to match it in sixteen others. Even when agents surpass human benchmarks, they do not reach the theoretical performance ceiling for the underlying tasks. These findings indicate that AIRS-Bench is far from saturated and offers substantial room for improvement. We open-source the AIRS-Bench task definitions and evaluation code to catalyze further development in autonomous scientific research.
[…]
1 Introduction
From the rise of deep learning through to the current era of Large Language models (LLMs), the rapid progress in machine learning (ML) has been largely benchmark-driven, making robust evaluations a key requirement (Hardt, 2025). This trend is reinforced by the current reviewing crisis, which highlights the limits of human judgment in assessing the quality of research contributions (Xu et al., 2022; Lawrence, 2022). Recent improvements in LLM technology have extended their abilities from simple tasks to complex agentic workflows, including scientific reasoning and coding (Khatri et al., 2025; Andrews et al., 2025; Lu et al., 2024; Yamada et al., 2025). Whilst early LLM research focused on improving the capabilities of LLMs through better data and pre-training, recent work has focused on extending their abilities by leveraging more test-time compute (Yao et al., 2023a), in particular by leveraging scaffolds that iteratively query the LLM at test-time so as to obtain a better response. Such scaffolds can enable a test-time search of the solution space by incorporating environment feedback (Nathani et al., 2025; Toledo et al., 2025) and yield much more performant solutions. Despite this potential, we lack a standardized framework to measure how well these agents perform the actual work of a research scientist. To fill this gap, we introduce AIRS-Bench.
We adopt the broadly accepted definition of an agent as a computer system situated in an environment that is capable of autonomous action in order to meet its design objectives (Wooldridge and Jennings, 1995). In our context, an agent is characterized as an LLM augmented by a scaffold. The LLM is the underlying probabilistic model that serves as the agent’s primary reasoning core, either self-hosted OSS model or API-based. The scaffold acts as an orchestrating layer, providing the necessary mechanisms to translate LLM outputs into systematic exploration of the solution space. The scaffold is typically implemented by a harness, the execution framework responsible for instantiating and managing diverse scaffolding configurations. Examples are the ReAct scaffold (Yao et al., 2023b), implemented within the MLGym harness (Nathani et al., 2025) or the Monte Carlo Tree Search (MCTS) scaffold (Kocsis and Szepesvari, 2006), implemented within the AIRA-dojo harness (Toledo et al., 2025). We experiment with both sequential (linear) and parallel scaffolds. Sequential scaffolds follow a linear execution loop, where each LLM query is conditioned on feedback from previous actions (Yao et al., 2023b; Nathani et al., 2025). Parallel scaffolds, by contrast, maintain and grow a population of potential solutions, utilizing data structures such as trees to guide exploration (Novikov et al., 2025; Sharma, 2025; Toledo et al., 2025).
A growing area of research for LLM agents is represented by the automation of AI research itself. We call agents designed to automate and accelerate AI research AI Research Agents. Accurately evaluating the performance of such agents, however, remains a significant challenge. ML has long struggled with reproducibility and statistical noise (Recht et al., 2018; Hardt, 2025), and the agentic paradigm adds further complications:
Data contamination: LLMs are trained on vast amounts of internet data, and they often memorize benchmark solutions. This makes it difficult to assess whether an agent is truly “reasoning” on a task.
Environmental standardization: agentic environments are difficult to standardize across different studies, and it is often unclear if success comes from the agent’s capabilities or the specific way the environment was built.
Computational cost: The high computational overhead of each autonomous run makes it difficult to conduct the extensive trials necessary to obtain statistically significant results.
These factors have exacerbated the evaluation crisis of AI Research Agents, as performance on existing benchmarks is increasingly obscured by pretraining leakage, inconsistent environment setups, and noisy empirical evaluations (Haimes et al., 2024; Dehghani et al., 2021).
We introduce the AI Research Science Benchmark (AIRS-Bench) to address these limitations and provide a standardized evaluation of AI Research Agents. The philosophy behind AIRS-Bench is that, much like the progress in ML, the advancement of AI Research Agents should be a benchmark-driven process. AIRS-Bench is a suite of 20 agentic tasks, curated from recent state-of-the-art (SOTA) literature. This ensures that tasks are both challenging and relevant to the broader ML community. In our design, agents are not required to tackle these tasks through direct inference. Instead, they must generate the code necessary to train and validate an ML model. We then evaluate the agent’s capabilities by executing the code it generated and measuring the model performance. This approach effectively evaluates the agent’s capacity to function as an autonomous research scientist.
To both standardize evaluations and make the tasks closely match their original versions in the ML literature, we developed a task configuration standard that encapsulates an AIRS-Bench task (see Figure 1 for an example) and a task creation pipeline based on semi-manual sourcing, creation, reviewing and verification of tasks. Our proposed AIRS-Bench task standard can be adapted and extended to virtually any ML problem, effectively democratizing AI agentic research. For fair evaluations across all agents we have designed our experiments to make runs as comparable as possible by factoring in infrastructure issues (e.g. crashing runs) and eliminating environment discrepancies. To investigate the impact of different scaffolds, we evaluate both the AIRA-dojo (Toledo et al., 2025) and MLGym (Nathani et al., 2025) harnesses combined with different LLMs (CWM (Carbonneaux et al., 2025), GPT-4o (OpenAI, 2024a), gpt-oss-20b and gpt-oss-120b (OpenAI, 2024b), o3-mini (OpenAI, 2024c), Devstral (Rastogi et al., 2025)). We also introduce an evaluation protocol, consisting of different metrics and aggregating results across seeds and tasks to allow for statistically robust and interpretable results.
The key contributions of the paper are:
Uncontaminated benchmark for AI Research Agents: AIRS-Bench uniquely assesses agents on the end-to-end ML research workflow: idea generation, methodology design, experiment analysis, and iterative refinement—without access to baseline code. This enables a realistic evaluation of agentic research abilities. AIRS-Bench tasks cover a variety of machine learning research problems, ranging from NLP, math and code to biochemical modeling and time series forecasting. The benchmark is designed to be compatible with multiple agentic frameworks (i.e. harnesses), supporting robust and fair comparisons across different agent architectures and tool integrations.
Task configuration standard: we introduce a task configuration standard and fixed evaluation metrics, to ensure reproducibility and minimize runtime and environment inconsistencies. We establish quality by leveraging human checks throughout task building, reviewing, and verification.
Empirical analysis: we benchmark all tasks across frontier open and closed-source models and different scaffolds, revealing substantial variation in performance. We closely inspect cases where agents are shown to match or beat human SOTA and analyze the agent-generated solutions—including cases where agents discover effective SOTA-exceeding combinations of approaches.
Evaluation protocol: we define a suite of metrics assessing different aspects of agents’ capabilities including valid submission rates, normalized performance scores, and Elo ratings. Normalization enables us to aggregate performance across tasks by applying transforms that map scores to the weakest valid solution and to human SOTA. The metrics used are designed to accurately reflect progress on developing state-of-the-art AI Research Agents. We report results on these metrics across the range of selected agents; our results suggest that the benchmark is far from solved, leaving plenty of headroom for development of AI Research Agents.
The rest of the paper is organized as follows. Section 2 contains an overview of existing benchmarks within the AI research domain. Section 3 introduces the agent, scaffold and harness definitions and section 4 presents the benchmark structure and creation methodology. Sections 5 and 6 outline our experimental design and results. We conclude with discussing learnings and key findings in Section 7.
2 Related Work
We examined a number of benchmarks evaluating the ability of AI agents to carry out scientific research, including ML-related agentic tasks. Below, we outline several trends we identified and how they relate with AIRS-Bench.
Full cycle of Scientific Method. Many earlier papers create agentic tasks from available sources: Github repositories (CSR-bench; Xiao et al., 2025), top-tier conference papers and their data (PaperBench; Starace et al., 2025a), Kaggle competitions (MLE-bench; Chan et al., , DSBench; Jing et al., 2025), ML research competitions (MLRC-Bench; Zhang et al., 2025), and carefully selected cross-domain papers (SciReplicate-bench; Xiang et al., 2025b). The experimental cycle of these benchmarks can include ideation, implementation, experimentation, analysis and comparison to previous results, with several benchmarks focusing on separate stages of the cycle, such as IdeaBench (Guo et al., 2024), LiveIdeaBench (Ruan et al., 2025), AI Idea Bench (Qiu et al., 2025) and ResearchBench (Liu et al., 2025b) for ideation, FML-bench (Zou et al., 2025) for idea novelty estimation, SurveyBench (Sun et al., 2025) for literature review, and DataGovBench (Liu et al., 2025c), DCA-Benchmark (Huang et al., 2025), DA-Code (Huang et al., 2024b), and DS-1000 (Lai et al., 2022) for data quality, implementation and analysis. AIRS-Bench, on the other hand, requires the agent to excel in every step of the scientific method to perform well on the benchmark tasks, spanning hypothesis generation, implementation, experimentation, and analysis.
| Benchmark | Task Composition and | AI Research | Task | Scientific Method | No | Agent | |||
| Origin | Data | Horizon | (H) | (I) | (E) | (A) | Baseline | Compute | |
| AIRS-Bench (ours) | 20 tasks from 17 machine learning | ||||||||
[…]
| Automated LLM | 76 tasks from the NanoGPT | ||||||||
[…]
| Auto-Bench | 6 graph discovery tasks | ||||||||
[…]
| CORE-Bench | 270 tasks from 90 social, medical, | ||||||||
[…]
| CSR-Bench | 107 Github repos from CV, NLP | ||||||||
| (Xiao et al., 2025) | and interdisciplinary ML papers | ✓ | not specified | x | ✓ | ✓ | ✓ | x | not specified |
| MLE-Bench | 75 Kaggle | ||||||||
[…]
| MLGym-Bench | 13 tasks from supervised learning, | ||||||||
[…]
| ML-Agent-Bench | 13 tasks from Kaggle, computer | ||||||||
[…]
| SWE-Bench | 2,294 Github issues from | ||||||||
[…]
| SciReplicate-Bench | 100 tasks from | ||||||||
[…]
| PaperBench | 8,316 rubrics from | ||||||||
[…]
| ResearchBench | Hypothesis generation task | ||||||||
| (Liu et al., 2025a) | for 1386 papers across the sciences | ✓ | not specified | ✓ | x | x | ✓ | ✓ | not specified |
| RE-Bench | 7 LLM pretraining/coding tasks | ||||||||
[…]
| LMR-Bench | 28 code reproduction tasks | ||||||||
| (Yan et al., 2025) | from 23 NLP papers | ✓ | not specified | x | ✓ | x | x | x | not specified |
| PostTrainBench | Post-train 4 base LLMs to | ||||||||
[…]
Access to Baseline Solution. Benchmarks differ significantly in two ways: whether the agent is granted access to a baseline solution, and the degree of “saturation” within the underlying problem. AIRS-Bench prioritizes unsaturated tasks, and does not provide a starter solution to the agent. This approach makes AIRS-Bench tasks challenging, as agents must navigate a longer reasoning horizon to make progress independently.
Different environments. Recent benchmarks leverage a wide range of different environments. While most setups closely resemble AIRS-Bench—requiring agents to interact via prompts and code execution—others incorporate gamified environments (DiscoveryWorld; Jansen et al., 2024) or physics engines (FEABench; Mudur et al., 2024). Task sets are typically manually curated, though some benchmarks (DiscoveryBench; Majumder et al., 2024, SUPER; Bogin et al., 2024) also include synthetic tasks. Some papers simulate user interactions within the environment (AppWorld; Trivedi et al., 2024).
Diversity of domains. Beyond core ML tasks, there is a growing trend to expand agentic evaluation into broader scientific domains, such as bioinformatics, chemistry, and physics. We identify more than 20 different scientific domains covered across recent benchmarks, including DiscoveryWorld (Jansen et al., 2024), DiscoveryBench (Majumder et al., 2024), CURIE (Cui et al., 2025), FEABench (Mudur et al., 2024), ScienceAgentBench (Chen et al., 2025b), CORE-bench (Siegel et al., 2024b), AUTObench (Chen et al., 2025a), ResearchBench (Liu et al., 2025b) and BioMLbench (Miller et al., 2025). AIRS-Bench features text and tabular-based tasks across seven categories of ML problems, including language modeling, mathematics, code generation, molecular modeling, and time-series forecasting. Currently, AIRS-Bench provides a comprehensive testbed for agents, with the potential for future expansion into additional scientific domains.
We selected a subset of benchmarks that most closely resemble AIRS-Bench—specifically those evaluating an agent’s ability to conduct AI research independently. Table 1 summarizes this comparison across several dimensions: task composition and origin, reasoning horizon, the stages of the scientific method covered, access to baseline solutions, and compute requirements. AIRS-Bench is composed of tasks grounded in AI research data and crafted so that AI research agents (i) tackle long-horizon research challenges that mirror the complexity of scientific discovery (ii) engage with the complete scientific pipeline from hypothesis generation to implementation, experimentation and analysis in an iterative fashion (iii) operate without a starting solution to ensure unbiased evaluation and promote original research and (iv) leverage substantial computational resources to enable thorough solution space exploration and tackle modern AI research problems. These characteristics make AIRS-Bench well-suited for tracking progress toward developing state-of-the-art AI Research Agents.
3 Agents, Scaffolds, Harnesses
In line with the broader agent research literature, we view an agent as a computer system that is situated in some environment and and is able to act autonomously in this environment in order to achieve its design objectives (Wooldridge and Jennings, 1995). For our specific context of LLM agents for AI Research, we define an agent as the combination of an LLM and a scaffold. A scaffold is an algorithm, expressed as a a set of operators and search policies governing how the agent searches the space of possible solutions for the task at hand. Scaffolds are instantiated within a harness—a system that wraps the agent and manages its execution (see Figure 2). Examples of harnesses include AIDE (weco.ai, 2024; Jiang et al., 2025), Claude Code, 111https://code.claude.com/docs/ OpenCode, 222https://github.com/anomalyco/opencode MLGym (Nathani et al., 2025) and AIRA-dojo (Toledo et al., 2025). For instance, we refer to greedy search as a scaffold that is instantiated within the AIRA-dojo harness.
AIRA-dojo is a harness that enables the agent to evolve its solution through a set of operators and a search policy. The search policy (e.g., greedy search, Monte Carlo Tree Search (Kocsis and Szepesvari, 2006), evolutionary algorithms (Novikov et al., 2025; Sharma, 2025; Romera-Paredes et al., 2023)) guides the exploration of the solution space, while the operators modify existing solutions to generate new candidate solutions. This process results in a tree structure, where each node represents a Python code solution, created by one of the following operators: (1) Draft, which generates the initial set of solutions; (2) Debug, which identifies and corrects errors within a given node; and (3) Improve, which enhances a solution to increase its performance according to specified evaluation criteria. AIRA-dojo operators enhance AIDE to (i) promote solution diversity, by inserting into the context solutions of sibling nodes, and (ii) facilitate bug fixing, by including in the context the entire ancestral memory of the solution’s debug chain. AIRA-dojo allows access to the internet, including additional data and pretrained models, which are loaded in the cache for the agent to use according to prompting instructions. AIRA-dojo’s operator prompts can be found in Appendix C.2.
MLGym is a harness through which an agent can sequentially improve its initial solution in a ReAct-like manner (Yao et al., 2023b). The agent can develop a solution based on its own ideation and feedback from the execution of the implementation. Access to the internet is also allowed in a manner similar to AIRA-dojo, with the option to cache pretrained models and instruct the agent about them in the prompt. The agent has access to tools and bash, and the full system prompt is present in Appendix C.1.
4 Method
AIRS-Bench consists of tasks extracted from machine learning papers and different datasets. To ensure selection of highly impactful tasks for the community, we sourced tasks and SOTA results from papers published at well-known AI conferences and journals and arXiv preprints (see Appendix F for a distribution of publication venues and years). The initial list of tasks was sourced from PapersWithCode’s leaderboards (Taylor, 2020; Kardas et al., 2020) and filtered down for datasets that had a state-of-the-art result published between 2020 and 2025; additional criteria applied at this stage included availability of the dataset (preferably via HuggingFace) and existence of a train and a test split. After verifying these requirements, we manually created each task and reviewed the SOTA result associated with it, by confirming the number reported in the cited paper, checking that the evaluation metric and dataset splits were identical and updating the SOTA result if a better score could be found. By following this methodology, we initially created and evaluated approximately 100 tasks from 85 different machine learning papers and datasets. We subsequently selected 20 out of these to ensure tractability and accuracy of the benchmark. The process is outlined in Appendix A.
4.1 Tasks Diversity
Figure 3 show the distinct categories that AIRS-Bench’s tasks are organized into, with the tasks spanning a broad spectrum of ML problems. The largest share is dedicated to the NLP domain, comprising Question Answering (4/20, open-ended and extractive question answering over diverse contexts), Text Extraction & Matching (3/20, e.g. coreference resolution, semantic similarity) and Text Classification (2/20, sentiment analysis, document categorization) tasks. Then, Molecules and Proteins ML tasks (5/20) focus on predicting molecular properties and solving biochemical modeling problems, whereas Time Series tasks (3/20) perform forecasting over temporal data. Finally, Code (2/20) and Math (1/20) tasks correspond to code generation/retrieval and mathematical reasoning, respectively.
4.2 Key Task Fields
Table 4.2 lists the core task fields of an AIRS-Bench task. Specifically, name contains the unique task name, research_problem describes the problem the agent is asked to solve, dataset contains the HuggingFace dataset identifier, config corresponds to the subset of the HuggingFace dataset used and train_split and test_split to the HuggingFace datasets splits used by the agent for training and evaluation, respectively. The metric field contains the metric the agent is asked to minimize or maximize (depending on the value of metric_lower_is_better), while input_columns contains the dataset columns the agent can utilize to solve the research problem and scoring_column is the column the agent’s solution will be evaluated against.
Finally, category describes the domain of the research problem, sota_paper_title, sota_paper_url, sota_year and sota_venue hold information related to the paper containing the state-of-the-art result and sota_score is the state-of-the-art value of the metric appearing in the paper; the agent is not accessing any of the information stored in the category and sota-related fields.
lX
Field Value
name MathQuestionAnsweringSVAMPAccuracy
research_problem Math Question Answering
dataset ChilleD/SVAMP
metric Accuracy
metric_lower_is_better false
config default
train_split train
test_split test
input_columns question_concat
scoring_column Answer
category Math
sota_score 0.942
sota_year 2026
sota_venue Frontiers of Computer Science
sota_paper_title Achieving >97% on GSM8K: Deeply Understanding the Problems Makes LLMs Better Solvers for Math Word Problems
sota_paper_url https://arxiv.org/pdf/2404.14963v5
dataset_paper_url https://arxiv.org/abs/2103.07191
4.3 Task Files
Table 4.2 describes most of the key information that forms an AIRS-Bench task. This information requires additional code and data to be run on the AIRA-dojo and MLGym scaffolds. Sections C.1 and C.2 of the Appendix contain the system prompts of the two scaffolds. The full task specification includes a folder, whose name is the same as the task name, and a number of files that are required for the agent to solve the task within AIRA-dojo. The task files are organized in such a way that they can be easily and programmatically converted into files required by different harnesses; we show this by converting them into task definition files for the MLGym agentic framework. The linked Github repository333https://github.com/facebookresearch/airs-bench contains the AIRS-Bench task specifications for AIRA-dojo and MLGym along with scripts for AIRA-dojo-to-MLGym format conversion and scripts for preparing the experimental environment (e.g. dataset downloads). In the remaining of this section, we describe the format and purpose of the task definition files for AIRA-dojo.
4.3.1 project_description.md
The project_description.md file contains the instructions provided to the agent to complete the task in the form of a lengthy and appropriately structured prompt; see Appendix B.1.1 for a full example. The prompt is divided into three sections: a description of the research problem, a description of the dataset and an explanation of the evaluation setup.
The research problem is presented in a sentence that describes the objective of the problem along with the column of the dataset that the predictions will be evaluated against. For the running example, this is:
“Your task is to solve math word problems. Each example presents a short story followed by a specific question. Your task is to read the text and predict the correct numerical answer. Your predictions will be scored against the Answer column of the test set
In the dataset description, we report the structure of the HuggingFace dataset. In particular, we specify which repo the data comes from and the dataset schema (features) with an overview of the columns and their datatypes. This helps the agent understand how the data looks like even if the scaffold does not provide a lookahead function. All the data used by the agent during a task is pre-downloaded and exported to the agent’s container.
Lastly, we explain to the agent how to submit the solution and how it will be scored: the agent is expected to submit its solution in the form of a .csv file containing the predictions on the test split, which has the benefit of a standard output format. We also provide the agent with the code of the evaluation script (evaluate.py, explained below) that contains the metric implementation and will be used to score the agent-produced submission.csv file against the test data.
4.3.2 prepare.py and evaluate_prepare.py
The prepare.py and evaluate_prepare.py files contain the one-time data preparation logic for the agent to solve the problem and for the evaluation of the agent’s solution, respectively. Please note that there are differences between the two settings, as the test labels need to be removed while the agent is building its solution; the two scripts take care of these requirements.
4.3.3 evaluate.py
The evaluate.py file is the evaluation script used to score the agent’s submissions against the test data. See Appendix B.1.2 for the evaluation script of the MathQuestionAnsweringSVAMPAccuracy task. The script contains three core functions: load_test_set where the script loads the test data, evaluate which implements the metric used to score the submissions, and cli which orchestrates loading of the agent’s submissions and test data, running the evaluate method on these, and reporting the results to stdout.
4.3.4 metadata.yaml
The metadata.yaml file contains all the metadata about the task (same as the fields of Table 4.2 described in detail above) along with additional requirements to run the task (like libraries used by the evaluation script that need to be installed). See Appendix B.1.3 for the metadata.yaml file of the MathQuestionAnsweringSVAMPAccuracy task.
4.3.5 utils.py
The utils.py file is an optional file to consolidate overlapping code between the prepare.py, evaluate.py and evaluate_prepare.py files. Examples include normalization transforms used for data preparation or bespoke label extraction logic.
4.3.6 Train and test datasets
All datasets have been downloaded beforehand using the prepare_hf_datasets_text.py script. The data required for the task is then mounted to the agent’s container and prepared using the code within prepare.py and evaluate_prepare.py. The data folder always contains a train split under a folder whose name is the value of the train_split field of metadata.yaml and a test split under a folder similarly named using the test_split field. For the train split, one preparation step is removing all but the relevant columns for the task (and transforming their content if needed). This is to ensure that the model does not have access to extra data that might hint to the solution. The test split contains the test labels when used by the evaluation script to score the agent’s submissions, but it does not contain the labels when accessed by the agent to look at the structure of the test set and solve the problem.
5 Experiments
5.1 Evaluation Setup
We evaluate AIRS-Bench using two harnesses, MLGym and AIRA-dojo. To isolate the effects of design choices associated with different harnesses, we ensured similar constraints and resources across all runs. For further details, refer to Table 7 in Appendix C. The greedy scaffold within AIRA-dojo explores several solutions through a tree-based search policy, while MLGym operates sequentially within one reasoning stream. Each run lasts for 24 hours with access to one H-200 GPU. We launch each run of each task at least 10 times (which we refer to as 10 “seeds”).
Throughout AIRS-Bench evaluations with MLGym and AIRA-dojo harnesses, agents are allowed to access HuggingFace checkpoints, permitting the use of pretrained models. To facilitate this process and to mitigate HuggingFace rate limits, we locally cache a number of pretrained checkpoints. Note that the cache does not offer access to the latest foundational models and the most recent cached model dates back to . The full list can be found in section E of the Appendix. For both AIRA-dojo and MLGym runs, agents were explicitly instructed about the existence of the cache. Across all experiments, we do not provide agents with any information regarding the methodology or score of the SOTA paper. We hypothesize that some of the tasks in the benchmark would have benefited from more compute or time, but we kept costraints uniform across tasks to provide the agents with the same resources and push the limits of their ideation capabilities.
5.2 Metrics and Score Aggregation
We evaluated the performance of the agents on AIRS-Bench using three metrics: mean valid submission rate, average normalized score and Elo rating. The definitions of these metrics are provided below. Throughout all metrics and empirical results, we follow the terminology introduced in Section 3, for which an agent is defined as the combination of a scaffold (e.g. Greedy) and a base LLM (e.g. gpt-oss).
For each task, the agents faced the challenge of being able to submit a valid solution, i.e. one that meets the requirements specified in the task description and yields a valid score. Our first evaluation metric is thus the mean valid submission rate (VSR) across tasks for an agent , defined as
where is the number of valid (successful) runs for agent on task , is the number of total runs for agent on task , is the number of tasks over which agent is being evaluated (i.e. the AIRS-Bench tasks) and agent is a combination of a base LLM with a scaffold. The mean valid submission rate assesses the agents’ capability to come up with a working solution and submit it confidently.
Producing an aggregate score for AIRS-Bench is challenging due to the high diversity of tasks included: most tasks have unique metrics, and even for tasks sharing the same metric (e.g. accuracy), ranges reported in the literature for each of them may vary significantly. To aggregate heterogeneous metrics and ranges into a common scoring system, we define the normalized score (NS) of an agent on a task as:
where corresponds to the worst score observed across all seeds and all agents on task , is the SOTA score on task sourced from literature, is the score achieved by agent on task and is a non-linear transformation. Note that and will always correspond to normalized scores of and , respectively. Equation 2 involves a two-step normalization: first, for a given task , we apply a monotonic map onto the raw score achieved by the agent’s submission; second, we restrict the resulting scores within the interval if the agent performs worse than SOTA and if the agent exceeds SOTA to enable subsequent aggregation across tasks. Specifically, we employ the march of 9s transform444The ”march of nines” is a metaphorical expression, popularized by AI researcher Andrej Karpathy, to describe the vast and non-linear engineering effort required to achieve higher levels of reliability in AI systems (Karpathy and Patel, 2025). as our choice of , defined as
where is the overall possible optimal score for the task (e.g., for classification accuracy, for regression error), as opposed to the best score obtained or SOTA (which e.g. for accuracy would be less than 1.0). This choice of is to adjust changes of the score so that they reflect intuitive measures of progress on the benchmark: this approach treats closing e.g. the gap from to as significant as closing the gap from to , since both represent a tenfold reduction in the distance to optimal.555The reader can contrast this with a simpler but less representative transform definition, such as the identity transform, for which we present results in Section B of the Appendix When averaging normalized scores across seeds, we include both failed (i.e., the agent fails to submit a valid solution) and invalid submissions (i.e., the agent submits a solution that does not yield a numerical score) treating them as submissions with 0 normalized score.
Lastly, to quantify the relative skill level of each agent evaluated, we employ the Elo rating system (Elo, 1967). We do so by treating each agent as a player. For each AIRS-Bench task, we treat each pairwise comparison of the agents’ scores as a game, with the agent producing a better score winning that game. If two agents do not produce a valid submission or both produce the same score, that game is considered a tie.
We estimate the agents’ ratings by fitting a Bradley–Terry (BT) model to the head-to-head outcomes, following the approach used in Chatbot Arena (Chiang et al., 2024). The model infers latent skill parameters ( for agent ) such that the probability of one agent outperforming another agent follows a logistic function of their skill difference:
We then convert the estimated skill parameters into Elo ratings using the following transformation, where N denotes the total number of evaluated agents:
Unlike the classical Elo calculation, the BT model is order-invariant, making it well-suited for batch evaluations across multiple agents and tasks. Please note that in our setting we treat the human SOTA scores as an additional agent, the “SOTA” agent.
6 Results
We evaluate a total of 14 agents, i.e. LLM-scaffold pairs. The language models used are the Code World Model (CWM), o3-mini, gpt-oss-20b and gpt-oss-120b, GPT-4o and Devstral-Small 24B. We evaluate three scaffolds: (i) One-Shot, where the agent can attempt solving the problem only once with the same set of AIRA-dojo operators (by definition that would be the Draft operator only) (ii) Greedy, where the agent performs greedy search with the AIRA-dojo operator set and (iii) ReAct, where the ReAct prompting technique implemented by MLGym is powering the agent.
6.1 Comparing performance across agents
Figure 4 provides an overview of the three aggregate benchmark metrics introduced above. We observe that reasoning models (e.g. gpt-oss-120b, o3-mini) perform better in both one-shot and greedy settings. However, model size seems to not affect performance with the use of test-time scaling making up for it, i.e. Greedy gpt-oss-120b is on par with Greedy gpt-oss-20b. Moreover, tree-search methods benefit agents powered by both open-source and closed-source models, as suggested by e.g. the sizeable gaps between the performances of Greedy CWM and One-Shot CWM as well as Greedy GPT-4o and One-Shot GPT-4o agents. At the same time, performance of agents backed by linear scaffolds, such as ReAct CWM and ReAct GPT4o varies; in one case the ReAct agent is better than its greedy counterpart (CWM) and in the other case worse (GPT4o). We observe that the relative ranking of agents shows similar (but not identical) trends for the average normalized score and Elo Rating performance metrics. However, the ability to submit a valid solution does not always correlate with the ability to submit a high-performing solution. This suggests that agents who are willing to take risks may achieve better performance. Finally, the majority of agent-task combinations achieve results between the worst possible score and human SOTA, with a small but notable fraction (1.58%) exceeding SOTA, primarily driven by greedy search strategies.
We present valid submission rates for each agent in Figure 5, highlighting different ranges of valid submission rate across all tasks. We consider four submission rate ranges: invalid indicates that the agent failed to submit a valid solution across all its seeds; low/medium/migh correspond to the 1–33%, 34–66% and 67–100% of valid submission rates. Agents Greedy gpt-oss-120b and Greedy gpt-oss-20b lead with the smallest fractions of tasks yielding an invalid submission, at 6% and 7% respectively.
A breakdown of each agent’s average normalized score per task is provided in Figure 6. For each agent, we report the percentage of tasks for which the agent yields one of five possible outcomes: (i) invalid, the agent does not submit a solution at all; (ii) worst, the agent produces the lowest score among all agents for that task; (iii) below average, the agents achieves a score below the average across all agents for that task; (iv) above average, the agent achieves a score above the average but is not the best; and (v) best, the agent achieves the highest score among all agents for that task. This breakdown highlights the distribution of each agent’s performance across the benchmark tasks. Scores are normalized according to Equation 2, where is the march of 9s transform.
We report the mean valid submission rate defined in Equation 1 across the AIRS-Bench tasks in Figure. 7. On average, only 55.1% of the total submissions are considered valid, suggesting that even submitting a valid solution stretches the capabilities of the agents. We also observe that reasoning models and the Greedy/ReAct scaffolds offer an advantage, as performance of these agents is superior.
In Figure 8 we report the average normalized scores according to Equations 2 and 3. A more detailed breakdown of the average scores per task is reported in Figure 9 with specified by Equation 3. The scores distribution in the figure reiterates the value of the AIRA-dojo harness in supporting agents to develop better solutions, with Greedy scaffolds (in red) distributing closer to SOTA than One-Shot ones (in blue).
Figure 9 depicts normalized scores computed according to Equations 2 and 3: each row corresponds to a task and each point relates to the normalized score achieved by an agent on that task and averaged across multiple seeds. For each task we average the normalized score across all agents and seeds and we use this mean normalized score to sort tasks by difficulty level. We stack tasks from the easiest to the most difficult going from top to bottom. The mapping between task numbers appearing in the y axis of Figure 9 and task names can be found in Table LABEL:tab:task-key of Appendix B. Tasks with points to the right of the line indicate that the average score of that agent on that task exceeds human SOTA, which is not necessarily the case for all tasks with at least one seed (i.e. agent run) exceeding human SOTA. Figure 12 in Appendix B similarly presents normalized scores across tasks and agents, but using in Equation 2 the identity transform from Equation 8 (in Appendix B) instead of the march of 9s transform from Equation 3.
Based on the task ranking from Figure 9, in Figure 10 we group the AIRS-Bench tasks into four groups, each containing 5 tasks, and corresponding to an increasing level of difficulty: easy (tasks 1 to 5), medium (tasks 6 to 10), hard (tasks 11 to 15) and expert (tasks 16 to 20). Here we are averaging scores across 5 tasks, i.e. each point is the average over seeds of all 5 tasks in that difficulty bucket. We also observe in Figure 10 that while the normalized scores are all low and somewhat similar on the expert tasks, for the easier problems (and especially those in the easiest bucket), we see high variability between the scores that the agents achieved. The correspondence between task numbers and names is the same as the one in Figure 9 and can be found in Table LABEL:tab:task-key of Appendix B.
Finally, Elo ratings including human SOTA as an additional opponent alongside our agents are reported in Figure 11. The sizeable gap between the rating of the human SOTA player and the top-performing agent indicates that even the best agent is significantly below SOTA and the benchmark is very far from saturated.
6.2 Task Inspection: Success in Beating SOTA
Among the runs shown in Figures 9 and 12, we found cases where the agent’s performance, at least in some of the seeds, was higher than the reported human SOTA, i.e. had a normalized score that was greater than . Overall, we identified tasks where our agents surpass SOTA performance, as summarized across Tables 3-5. We examined these cases in depth to better understand the solution produced by the agent, and how it manages to outperform the human SOTA. Below, we provide a breakdown for a notable case where an agent outperforms human SOTA with an original solution.
Greedy gpt-oss-120b on TextualClassificationSickAccuracy. This is a Natural Language Inference (NLI) problem, and it employs the SICK (Sentences Involving Compositional Knowledge) dataset (Marelli et al., 2014a). Given a pair of sentences, a premise and a hypothesis, the goal is to determine the relationship between them, including: entailment (the hypothesis is true given the premise), contradiction (the hypothesis is false given the premise); and neutral (no conclusion can be drawn on the hypothesis given the premise). The evaluation metric is accuracy.
The SOTA solution (Kalouli et al., 2023) is achieved by fine-tuning RoBERTa (Liu et al., 2019) on the original SICK training set and testing on the original SICK test set. The approach is straightforward: a single transformer model is fine-tuned on a specific training set, yielding a test accuracy of .
The Greedy gpt-oss-120b agent, on the other hand, produces a two-level stacked ensemble that combines multiple transformer models and a meta-learner. RoBERTa-large and DeBERTa-v3-large (He et al., 2023), are independently fine-tuned on the SICK training set. Each model processes sentence pairs and outputs logits for each class. The training is performed using 5-fold stratified cross-validation, ensuring robust out-of-fold (OOF) predictions and preventing overfitting. The logits from both base models are concatenated to form a feature vector for each example. These stacked features are then used to train a logistic regression meta-learner, which learns to optimally combine the predictions of the two base models. During cross-validation, the meta-learner is trained on OOF logits, ensuring that the meta-features are unbiased and not overfitted. At test time, the base models are retrained on each fold, their test logits are averaged, and the meta-learner uses these combined logits to make the final prediction. This architecture leverages the complementary information captured by RoBERTa and DeBERTa, and the meta-learner can exploit patterns that may not be apparent to either base model alone, achieving a test accuracy of .
[…]
| TextualClassificationSick- Accuracy | • SOTA: 0.90 • Agent: 0.93 | • SOTA (Kalouli et al., 2023): Vanilla fine-tuning of RoBERTa. • Agent: Finetuned RoBERTa-large and DeBERTa-v3-large base models using stratified cross-validation. Out-of-fold logits from both models employed to train a logistic-regression meta-learner that combines base models’ logits. Finally, base models retrained on all folds and meta-learner was used to produce final predictions. |
| TextualSimilaritySick- SpearmanCorrelation | • SOTA: 0.85 • Agent: 0.89 | • SOTA (Huang et al., 2024a): Finetuned RoBERTa-large and novel loss function (CoSENT). • Agent: RoBERTa-base and RoBERTa-large finetuned to predict similarity scores. Produces similarity scores using cosine similarity of frozen Sentence-BERT. Used cross-validation to learn weights for averaging the similarity scores produced by all three models. |
[…]
| CoreferenceResolution- WinograndeAccuracy | • SOTA: 0.85 • Agent: 0.88 | • SOTA (Lin et al., 2020): Fine-tune T5-3B in a text-to-text setup, scoring answer options by output token probabilities (“entailment” vs. “contradiction”). • Agent: Vanilla finetuning of DeBERTa-v3-large with classifier head. |
[…]
| TimeSeriesForecasting- RideshareMAE | • SOTA: 1.185 • Agent: 1.153 | • SOTA (Gong et al., 2025): General transformer-based time-series foundation model (not finetuned on this dataset). • Agent: Trains a Bi-directional GRU |
7 Conclusion
We introduced AIRS-Bench, a benchmark designed to rigorously evaluate the autonomous research capabilities of LLM agents in machine learning. Our benchmark covers 20 diverse, non-contaminated tasks spanning multiple domains, and is specifically constructed to assess agents across the full research workflow—from ideation and methodology design to experimentation and iterative refinement—without access to baseline code. Our results indicate a high variability in task performance, depending on both the LLM that the agent uses, and the harness it is based on. For most tasks, even the best performing agent is still significantly behind the human SOTA, showing that the benchmark is far from saturated. For a few tasks, our top agents managed to identify a solution outperforming the human SOTA. It would be interesting to see how much AI research agents can push the state-of-the-art further.666Note that even in the cases where the agent outperforms human SOTA, it is still interesting to see how far ahead the agent’s solution is over the human baseline. Hence, tasks where the agent normalized score exceeds 1.0 are still interesting benchmark tasks.
We gathered a number of useful takeaways during building of AIRS-Bench and evaluating a number of agents across its tasks:
Gaps in community infrastructure: within the current state of AI research, the task of tracking up-to-date SOTA became more challenging than ever. Both the growing amounts of paper submissions,777https://forum.cspaper.org/topic/76/submission-tsunami-at-neurips-2025-is-peer-review-about-to-collapse/2 the high compute cost of reproducing experiments on large models and the lack of unified platforms to represent results contribute to the situation. A new shared space with standardized format, updates, and machine-readable configurations for all published machine learning research is needed.
Performance gaps: several main factors cause the performance of the agents to be lower than it could be. The combinations of base LLMs with different scaffolds can lead to 1) problems with formatting, specifically submitting the correct solution after the experiments, 2) problems with saving intermediate results, 3) performance deterioration on main capabilities due to the context overflow. Longer agentic traces also lead to increased probabilities of misaligned behaviours and accumulated issues around code edits and debugging, which make it difficult to adhere to the methodology of the ML experimentation.
Human bottlenecks: for the work to be scaled across new domains and a bigger volume of tasks, automatic task onboarding pipelines will be needed. Current human validation procedures prevent the expansion at scale.
Role of restrictions: The ML benchmarks for AI Agents tend to be computationally costly, both on the training and inference side. Given the significant resource constraints faced in agentic evaluations—such as computational costs, time limits, and token usage—we acknowledge the possible role of restrictions in the obtained results. Benchmark methodology commonly faces the choice of either evaluating the systems in the very well-defined restricted conditions or lifting most of them and comparing the best obtained results. Although we adhere to the first choice for the sake of future extensive ablations, lifting certain restrictions could enable more flexible and efficient agent behaviors in the future.
Our results demonstrate that scaffold design significantly impacts agent solution quality, highlighting opportunities to improve performance through algorithms that better leverage test-time compute. We release AIRS-Bench to help identify performance gaps in AI research agents and catalyze the development of better methods for accelerating scientific progress. As agentic capabilities advance, continued benchmark development will be essential. We hope this benchmark fosters transparency, reproducibility, and rigorous, standardized evaluation of LLM agents in frontier research contexts.
Appendix
Appendix A Task Selection
We constructed AIRS-Bench by downsampling a pool of approximately 100 tasks to a representative subset of 20 tasks. The reduction to 20 tasks was implemented to substantially decrease GPU requirements and enable faster benchmarking. The AIRS-Bench subset was selected to closely mirror the full pool according to three key criteria:
Agent performance: each agent’s average score on AIRS-Bench is as close as possible to their average score on the full benchmark.
Category distribution: the proportion of tasks from each of the 7 categories in AIRS-Bench is as close as possible to that of the full pool.
Relative ranking fidelity: the ranking of agents by performance is preserved between AIRS-Bench and the full pool.
For each agent , we compute the mean normalized score over and over AIRS-Bench, where
where the normalized score is defined in Equation 2 (see Section 5.2) and is the performance score of agent on task . The AIRS-Bench subset is selected to minimize the mean absolute error (MAE) between and across all agents:
where is the set of all agents. To ensure representative coverage, is partitioned into four difficulty bands (, , , ), based on their relative ranking by average normalised score (see Section 6.1) and each containing approximately 25 tasks. AIRS-Bench is constructed by sampling a fixed number of tasks from each band. Four candidate difficulty band distributions were evaluated:
Uniform: tasks each from easy, medium, hard, and expert bands ().
Medium-skewed: easy, medium, hard, and expert tasks ().
Center-skewed: easy, medium, hard, and expert tasks ().
Medium-heavy: easy, medium, hard, and expert tasks ().
The final band allocation was selected by choosing the configuration that minimized the mean absolute error (MAE) between agent scores on the subset and the full benchmark. The search for the optimal subset was performed using three subset selection algorithms:
Random search: samples ten thousand candidate subsets, each respecting the band constraints, and retains the subset with the lowest MAE.
Simulated annealing: iteratively swaps tasks within bands, accepting both improvements and, with decreasing probability, worse solutions to escape local minima.
Genetic algorithm: evolves a population of candidate subsets through tournament selection, single-point crossover (), and mutation () minimizing MAE over generations.
Across all 12 possible ({algorithm} {band distribution}) combinations, the best-performing configuration was obtained using the genetic algorithm with a medium-skewed band allocation, achieving a minimum MAE of . Other competitive configurations included both uniform and skewed band distributions, with MAE values ranging from to .
To validate the fidelity of AIRS-Bench, we compared agent mean normalized scores and their 95% confidence intervals on AIRS-Bench versus the full pool . The results demonstrate that agent rankings and score gaps are faithfully preserved, with overall mean scores and confidence intervals nearly identical between the two sets: the difference in average score between the subset and the full benchmark never exceeds 0.02 in absolute value. This confirms that the selection criterion and stratified sampling approach yield a lightweight benchmark that maintains the discriminative power and ranking structure of the original pool, while allowing efficient evaluation.
Appendix B Additional Results
We consider an additional normalized score (see Eq. 2) which employs the identity transform
In contrast to Eq. 3, this approach directly uses raw scores. With this transform, the normalized score linearly reflects the agent’s progress between the worst observed solution and the human SOTA for each task. This is simple and interpretable, but may not always reflect meaningful progress when the metric is highly non-linear or when the gap to the optimal score is very small. For completeness, we show the average normalized score using this transform in Figure 12, and provide a breakdown of the scores by difficulty level in Figure 13.
For better readability, Table LABEL:tab:task-key provides the mapping between task numbers shown on the y-axis of Figures 9-12 and their corresponding task names, as well as their average score across all seeds and agents.
| Rank | Task (March of 9’s) | Task (Identity) | ||
[…]
B.1 Task run files
B.1.1 project_description.md
B.1.2 evaluate.py
B.1.3 metadata.yaml
Appendix C Harness Setup
| AIRA-dojo | MLGym | |
| Time/number of steps limit | 24 hours (up to 36 including evaluation time) | 1M steps / 24 hours |
| Can exit early | No | No |
| evaluate.py file is visible | Yes | Yes |
| Test set with labels is visible | No | No |
| Validation script | Agent codes during run | Agent codes during run |
| Validation splits | Cross-validation (Greedy only) | Classical split 70-30 |
| Scaffold implemented | Greedy (AIDE) | ReAct |
| Last submission valid always? | No | No |
| Pretrained models access | Yes | Yes |
| Num of steps / nodes captured | Yes | Yes |
| All validation scores captured | Yes | Yes |
| Every solution scored on test set | Yes | Only final submitted solution |
| Time limit per solution | 4 hours | 1 hour |
| Dummy submission provided | No | No |
| Num GPUs | 1 H200 / run | 1 H200 / run |
| Internet access | Yes (HF_OFFLINE=True, but agent can set it to False) | Yes (can be turned off) |
| How is evaluate.py provided | In prompt | In shared workspace |
| Python version | 3.10 | 3.10 |
| Datasets library version | 3.5.1 (upgrade to 4.0.0) | 4.0.0 |
C.1 MLGym system prompt
C.2 AIRA-dojo system prompt
The AIRA-dojo scaffold leverages the following predefined operator set: Draft initializes the search process by generating an initial population of candidate solutions. Debug attempts to identify and correct errors in buggy solutions. Improve refines valid artifacts to enhance their performance according to the evaluation criteria. While not explicitly defined as an operator, Analyze is also used to evaluate the execution output of a generated and executed solution to detect bugs and summarize empirical findings from the results. Detailed below are the prompts for each operator.
C.2.1 Draft
C.2.2 Debug
C.2.3 Improve
C.2.4 Analyze
Appendix D Compute Requirements of Benchmarks
p2.8cmX X X X X X
Benchmark GPU / Hardware Runtime Budget / Cost
Notes
AIRS-Bench 1×H200 GPUs per run 24h/task not specified
20 tasks in total, 10–20 runs/task
MLE-Bench 1×A10 GPU 24h/competition 1,800 GPU hours total
75 competitions in total
MLGym-Bench 0–2 GPUs per task (depending on task) 2–4h/task $1/run for most LLMs, some are up to $9
13 tasks in total
RE-Bench 0–6 H100 GPUs (depending on task) 8h/run $123/run
7 tasks in total; 3–5 runs/task
ML-Agent-Bench not specified 0.5–2h/task; max 5h/run $60 total
13 tasks in total
SWE-Bench not specified not specified $0.3 per task; $500 total
2294 tasks in total
CORE-Bench 1×T4 GPU or CPU 2h/task $4 per task; max $6; $500 total
270 tasks in total; 3 trials/task
CSR-Bench not specified not specified not specified
100 GitHub repos in total
Auto-Bench not specified not specified $365 total
6 tasks in total; 10-20 trials per task
SciReplicate-Bench 1×A100 GPU not specified not specified
36 papers broken down to 100 tasks; 3 runs / task
PaperBench 1×A10 GPU Up to 12h run/paper $400 / paper run; $8k total
20 research papers broken down to 8316 small tasks, 3 runs / paper
ResearchBench not specified not specified not specified
1386 papers each with 3 tasks
Automated LLM Speedrunning 8×H100 GPUs 10h/run; max 20h/run not specified
19 tasks each with 4 different levels; 3 runs per task+level
Appendix E Cached Models
The models’ cache available to our agents during the runs consists of the following pretrained models available on HuggingFace, as shown in Table LABEL:tab:allowed_models. This cache does not contain frontier models, the newest model present is deberta-v3-large released in .
| Model | Model |
| ai-forever–ruT5-base | ai4bharat–IndicBERTv2-MLM-only |
| ai4bharat–indic-bert | albert–albert-base-v2 |
| albert-base-v2 | albert-xxlarge-v1 |
| albert-xxlarge-v2 | allenai–longformer-base-4096 |
| allenai–scibert_scivocab_uncased | allenai–specter |
| anferico–bert-for-patents | BAAI–bge-large-en-v1.5 |
| BAAI–bge-small-en-v1.5 | bert-base-cased |
| bert-base-multilingual-cased | bert-base-multilingual-uncased |
| bert-base-uncased | bert-large-cased |
| bert-large-uncased | bert-large-uncased-whole-word-masking |
| bert-large-uncased-whole-word-masking-finetuned-squad | bhadresh-savani–bert-base-uncased-emotion |
| bhadresh-savani–distilbert-base-uncased-emotion | bhadresh-savani–roberta-base-emotion |
| camembert–camembert-base | camembert-base |
| cardiffnlp–twitter-roberta-base | cardiffnlp–twitter-roberta-base-emotion |
| cardiffnlp–twitter-roberta-base-sentiment | cardiffnlp–twitter-roberta-base-sentiment-latest |
| cointegrated–rut5-base | cointegrated–rut5-small |
| cross-encoder–ms-marco-MiniLM-L-6-v2 | cross-encoder–nli-deberta-v3-base |
| cross-encoder–nli-deberta-v3-large | cross-encoder–nli-roberta-base |
| cross-encoder–stsb-roberta-base | cross-encoder–stsb-roberta-large |
| deepset–roberta-base-squad2 | deepset–roberta-large-squad2 |
| deepset–xlm-roberta-base-squad2 | deepset–xlm-roberta-large-squad2 |
| DeepPavlov–rubert-base-cased | distilbert–distilbert-base-cased |
| distilbert–distilbert-base-uncased | distilbert–distilroberta-base |
| distilbert-base-cased | distilbert-base-cased-distilled-squad |
| distilbert-base-multilingual-cased | distilbert-base-uncased |
| distilbert-base-uncased-distilled-squad | distilbert-base-uncased-finetuned-sst-2-english |
| distilroberta-base | distilgpt2 |
| dmis-lab–biobert-v1.1 | facebook–bart-base |
| facebook–bart-large | facebook–bart-large-cnn |
| facebook–bart-large-mnli | facebook–fasttext-en-vectors |
| facebook–mbart-large-50 | facebook–mbart-large-cc25 |
| facebook–wav2vec2-base | facebook–wav2vec2-base-960h |
| facebook–wav2vec2-large-960h | FacebookAI–roberta-base |
| FacebookAI–roberta-large | FacebookAI–xlm-roberta-base |
| FacebookAI–xlm-roberta-large | gpt2 |
| gpt2-large | gpt2-medium |
| google–bigbird-roberta-base | google–bigbird-roberta-large |
| google–byt5-base | google–byt5-small |
| google–electra-base-discriminator | google–electra-large-discriminator |
| google–electra-large-generator | google–electra-small-discriminator |
| google–efficientnet-b0 | google–efficientnet-b3 |
| google–efficientnet-b4 | google–efficientnet-b5 |
| google–efficientnet-b6 | google–efficientnet-b7 |
| google–flan-t5-base | google–mobilebert-uncased |
| google–mt5-base | google–mt5-small |
| google–muril-base-cased | google–muril-large-cased |
| google–pegasus-xsum | google–t5-v1_1-small |
| google–vit-base-patch16-224 | google–vit-base-patch16-224-in21k |
| google–vit-large-patch16-384 | google-bert–bert-base-cased |
| google-bert–bert-base-multilingual-cased | google-bert–bert-base-uncased |
| google-bert–bert-large-uncased | google-bert–bert-large-uncased-whole-word-masking-finetuned-squad |
| google–bert_uncased_L-2_H-128_A-2 | google–bert_uncased_L-4_H-512_A-8 |
| google-t5–t5-base | google-t5–t5-small |
| helsinki-nlp–opus-mt-de-en | Helsinki-NLP–opus-mt-en-de |
| Helsinki-NLP–opus-mt-en-es | Helsinki-NLP–opus-mt-en-fr |
| Helsinki-NLP–opus-mt-en-ROMANCE | Helsinki-NLP–opus-mt-es-en |
| Helsinki-NLP–opus-mt-fr-en | Helsinki-NLP–opus-mt-ROMANCE-en |
| Helsinki-NLP–opus-mt-ru-en | j-hartmann–emotion-english-distilroberta-base |
| j-hartmann–emotion-english-roberta-large | joeddav–distilbert-base-uncased-go-emotions-student |
| llm-blender–PairRM | microsoft–codebert-base |
| microsoft–codebert-base-mlm | microsoft–deberta-base |
| microsoft–deberta-large | microsoft–deberta-v2-xlarge |
| microsoft–deberta-v2-xxlarge | microsoft–deberta-v3-base |
| microsoft–deberta-v3-large | microsoft–deberta-v3-small |
| microsoft–DialoGPT-medium | microsoft–graphcodebert-base |
| microsoft–MiniLM-L12-H384-uncased | microsoft–mpnet-base |
| microsoft–swin-base-patch4-window7-224 | microsoft–trocr-base-printed |
| microsoft–unixcoder-base | microsoft–xtremedistil-l6-h256-uncased |
| microsoft–xtremedistil-l6-h384-uncased | openai–clip-vit-base-patch16 |
| openai–clip-vit-base-patch32 | OpenAssistant–reward-model-deberta-v3-base |
| OpenAssistant–reward-model-deberta-v3-large | OpenAssistant–reward-model-deberta-v3-large-v2 |
| prajjwal1–bert-mini | prajjwal1–bert-tiny |
| princeton-nlp–unsup-simcse-roberta-base | ProsusAI–finbert |
| roberta-base | roberta-base-openai-detector |
| roberta-large | roberta-large-mnli |
| s-nlp–roberta_toxicity_classifier | Salesforce–codegen-350M-mono |
| Salesforce–codet5-base | Salesforce–codet5-base-multi-sum |
| Salesforce–codet5-large | Salesforce–codet5-small |
| SamLowe–roberta-base-go_emotions | sberbank-ai–ruT5-base |
| sberbank-ai–ruT5-large | sentence-transformers–all-distilroberta-v1 |
| sentence-transformers–all-MiniLM-L6-v2 | sentence-transformers–all-mpnet-base-v2 |
| sentence-transformers–msmarco-distilbert-base-tas-b | sentence-transformers–paraphrase-albert-small-v2 |
| sentence-transformers–paraphrase-MiniLM-L3-v2 | sentence-transformers–paraphrase-MiniLM-L6-v2 |
| sentence-transformers–paraphrase-mpnet-base-v2 | sentence-transformers–stsb-mpnet-base-v2 |
| siebert–sentiment-roberta-large-english | stanfordnlp–glove |
| t5-base | t5-large |
| t5-small | timm–efficientnet_b4.ra2_in1k |
| unitary–toxic-bert | unitary–unbiased-toxic-roberta |
| UrukHan–t5-russian-spell | vectara–hallucination_evaluation_model |
| vinai–bertweet-base | vinai–bertweet-large |
| xlnet–xlnet-base-cased | xlnet–xlnet-large-cased |
| xlnet-base-cased | xlm-roberta-base |
| xlm-roberta-large |
Appendix F Distribution of tasks SOTA venue and year
In Fig. 14 we reprot the breakdown of AIRS-Bench tasks by (a) SOTA publication venue and (b) SOTA publication year. A detailed breakdown of each venue is provided in Table 10.
| Venue | Count |
| ICLR | 5 |
| Preprint | 4 |
| ACL | 2 |
| AIMLSystems | 1 |
| Frontiers of Computer Science | 1 |
| Nature Communications | 1 |
| NEURIPS | 1 |
| ICML | 1 |
| EMNLP | 1 |
| Computational Linguistics | 1 |
| Model technical report | 1 |
| IEEE/ACM Transactions on Audio, Speech and Language Processing | 1 |
Appendix G AIRS-Bench Task Description
G.1 CodeGenerationAPPSPassAt5
Solve coding problems by generating five distinct Python programs for each problem. It employs the APPS dataset (Hendrycks et al., 2021), which consists of thousands of real-world coding challenges collected from online platforms, each accompanied by a detailed natural-language problem statement and a starter code template. For each test problem, the problem statement and starter code are provided. Each program is evaluated against a set of hidden test cases, and a prediction is considered correct if at least one of the five submitted programs passes all official test cases. Model performance is assessed using the Pass@5 metric, which measures the fraction of problems solved by at least one of the five attempts.
G.2 CodeRetrievalCodeXGlueMRR
Retrieve relevant code snippets given natural language queries. It uses the CodeXGlue Code Search Adv dataset (Lu et al., 2021), which consists of a large corpus of code functions in Java and a set of queries describing desired functionality in natural language. For each query, the task is to search the corpus and rank code snippets by relevance, aiming to identify the correct code that implements the described functionality. During training and validation, queries are paired with the correct code snippet, while in the test set, only the queries and the code corpus are provided. Model performance is assessed using the Mean Reciprocal Rank (MRR), which measures how highly the correct code is ranked for each query.
G.3 CoreferenceResolutionSuperGLUEWSCAccuracy
Predict whether a pronoun refers or not to something mentioned earlier in the sentence. It uses the SuperGLUE WSC dataset (Wang et al., 2019), which provides sentences with an ambiguous pronoun and a highlighted possible reference. For each example, the agent is given the sentence, the pronoun, and the possible reference. The goal is to predict whether the pronoun refers to that reference (binary classification). Model performance is measured using accuracy, which is the percentage of examples where the correct prediction is made.
G.4 CoreferenceResolutionWinograndeAccuracy
Identify which of two possible options a pronoun in a sentence refers to. It uses the Winogrande dataset (Sakaguchi et al., 2021), which contains sentences with an ambiguous pronoun and two possible answers. For each sentence, select the option that best fills the gap using commonsense reasoning. Model performance is assessed using accuracy, which is the percentage of correct answers.
G.5 CvMolecularPropertyPredictionQm9MeanAbsoluteError
Estimate a molecular property, the heat capacity at constant volume (), from the geometry and atomic composition of a molecule. The agent is required to perform regression based on the 3D coordinates of each atom and its corresponding element. The task utilizes the QM9 dataset (Ramakrishnan et al., 2014), a classic benchmark for molecular property prediction, which spans more than 10 different molecular properties determined using ab-initio density functional theory. Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values.
G.6 GMolecularPropertyPredictionQm9MeanAbsoluteError
Estimate a molecular property, the Gibbs free energy at (), from the geometry and atomic composition of a molecule. The agent is required to perform regression based on the 3D coordinates of each atom and its corresponding element. The task utilizes the QM9 dataset (Ramakrishnan et al., 2014). Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values.
G.7 GraphRegressionZincMae
Estimate a molecular property, the constrained solubility of a molecule, from its graph structure. The task utilizes the ZINC dataset (Sterling and Irwin, 2015), a widely used benchmark for graph-based molecular property prediction, which contains thousands of molecular graphs with associated solubility values. The agent is required to perform regression based on the molecular graph, where each molecule is represented as a graph with node features (atom attributes), edge indices (connectivity), and edge attributes (bond types). Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth solubility values.
G.8 MathQuestionAnsweringSVAMPAccuracy
Solve math word problems by reading a short story and answering a specific numerical question. The agent is required to predict the correct numerical answer based on the provided narrative and question, which may involve operations such as addition, subtraction, multiplication, or division. The task utilizes the SVAMP dataset (Patel et al., 2021), a benchmark for evaluating mathematical reasoning and problem-solving abilities in natural language. Each example consists of a description, a question, and the correct answer. Model performance is assessed using accuracy, which is the percentage of examples where the predicted answer matches ground-truth.
G.9 QuestionAnsweringDuoRCAccuracy
Answer questions based on a large context from movie plots. For each example, the agent is provided with the title of a story, a detailed plot summary, and a question about the story. The task is to determine whether the answer to the question is present in the context, and if so, to select the correct answer from a list of candidate answers. The DuoRC dataset (Saha et al., 2018) is used for this task, which contains diverse and challenging reading comprehension questions requiring reasoning over long narrative texts. Model performance is assessed using accuracy, which measures the percentage of questions for which the agent correctly identifies whether an answer exists and, if so, selects the exact answer from the provided candidates.
G.10 QuestionAnsweringEli5Rouge1
Answer open-ended questions using long-form, explanatory responses. For each example, the agent is provided with a question, a detailed context, and is required to generate a comprehensive, human-readable answer. The task utilizes the the ELI5 (Explain Like I’m Five) dataset, containing questions and high-quality, crowd-sourced answers (Fan et al., 2019). Model performance is assessed using the ROUGE-1 F-measure, which evaluates the overlap of unigrams (words) between the generated answer and the reference answer, measuring the quality and relevance of the response.
G.11 QuestionAnsweringFinqaAccuracy
Answer financial reasoning questions based on a combination of textual context and tabular data. For each example, the agent is provided with a question, supporting context, and a table containing relevant financial information. The task utilizes the FinQA dataset (Chen et al., 2021), which is designed to evaluate complex question answering and reasoning over both natural language and structured tables in the financial domain. Model performance is assessed using accuracy, which measures the percentage of questions for which the predicted answer exactly matches the ground-truth, accounting for both numerical and textual equivalence.
G.12 ReadingComprehensionSquadExactMatch
Extract answers to questions from context paragraphs in a reading comprehension setting. For each example, the agent is provided with a title, a context paragraph, and a question about the context. The task is to extract a span of text from the context that answers the question. The dataset uses the SQuAD dataset (Rajpurkar et al., 2016), which is a widely adopted benchmark for machine reading comprehension. Model performance is assessed using the Exact Match metric, which measures the percentage of predictions that exactly match one of the ground-truth answers provided in the dataset.
G.13 R2AbsMolecularPropertyPredictionQm9MeanAbsoluteError
Estimate a molecular property, the electronic spatial extent (), from the geometry and atomic composition of a molecule. The agent is required to perform regression based on the 3D coordinates of each atom and its corresponding element. The task utilizes the QM9 dataset (Ramakrishnan et al., 2014). Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values.
G.14 SentimentAnalysisYelpReviewFullAccuracy
Perform sentiment analysis on user-generated reviews from Yelp. For each example, the agent is provided with the text of a Yelp review and is required to predict the corresponding sentiment label, which represents the star rating assigned by the user. The dataset employed derives from the Yelp Dataset Challenge 2015 Asghar (2016), containing reviews labeled as one of five classes: ‘1 star’, ‘2 stars’, ‘3 stars’, ‘4 stars’, or ‘5 stars’ (encoded as 0, 1, 2, 3, or 4). Model performance is assessed using accuracy, which measures the percentage of reviews for which the predicted label exactly matches the ground-truth rating.
G.15 TextualClassificationSickAccuracy
Classify the entailment relationship between two sentences. For each example, the agent is provided with a pair of sentences A and B, and must predict whether the relationship is: entailment, i.e. sentence B can be logically inferred from sentence A; neutral, there is no clear logical relationship; contradiction, sentence B contradicts sentence A. The task uses the SICK dataset Marelli et al. (2014b), a standard benchmark for evaluating models on sentence-level semantic relatedness. Model performance is assessed using accuracy, which measures the percentage of predictions that exactly match the ground-truth label.
G.16 TextualSimilaritySickSpearmanCorrelation
Estimate the semantic relatedness between two sentences by predicting a similarity score from 0 (completely unrelated) to 5 (highly related). For each example, the agent is provided with a pair of sentences, and must output a floating-point score reflecting their degree of semantic similarity. The task uses the SICK dataset Marelli et al. (2014b). Model performance is assessed using the Spearman correlation coefficient between the predicted scores and the ground-truth scores, measuring how well the model’s ranking of sentence pairs matches the human-annotated rankings.
G.17 TimeSeriesForecastingKaggleWebTrafficMASE
Perform time series forecasting over the Kaggle Web Traffic dataset, which is part of the Monash Time Series Forecasting Repository. The repository is an extensive collection of time series datasets curated by Monash University and a widely adopted benchmark in the field (Godahewa et al., 2021). The dataset contains 145063 daily time series representing the number of hits or web traffic for a set of Wikipedia pages from 01/07/2015 to 10/09/2017 used by the Kaggle web traffic forecasting competition (Maggie et al., 2017). The goal of the task is to predict the future trajectory of the series by forecasting 59 time steps ahead. Model performance is assessed using the mean absolute scaled error (MASE) between the predicted and ground-truth values in the time series.
G.18 TimeSeriesForecastingRideshareMAE
Perform time series forecasting over the Rideshare dataset, which is part of the Monash Time Series Forecasting Repository (Godahewa et al., 2021). The dataset contains hourly time series representations of attributes related to Uber and Lyft rideshare services for various locations in New York between 26/11/2018 and 18/12/2018. The dataset contains 2304 individual time series, each capturing different aspects of rideshare demand and pricing, including pickup requests, pricing variations, and service availability across different geographic zones and time periods. The goal of the task is to predict the future trajectory of the series by forecasting 48 time steps ahead. Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values in the time series.
G.19 TimeSeriesForecastingSolarWeeklyMAE
Perform time series forecasting over the Rideshare dataset, which is part of the Monash Time Series Forecasting Repository (Godahewa et al., 2021). The dataset provides weekly aggregated solar power generation and forecast data for a large set of simulated photovoltaic (PV) plants across the United States. The dataset captures the dynamics of solar power generation, including seasonal variations, weather-dependent fluctuations, and geographic diversity across different climate zones. The goal of the task is to predict the future trajectory of the series by forecasting 5 time steps ahead. Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values in the time series.
G.20 U0MolecularPropertyPredictionQm9MeanAbsoluteError
Estimate a molecular property, the atomization energy at 0 K (), from the geometry and atomic composition of a molecule. The agent is required to perform regression based on the 3D coordinates of each atom and its corresponding element. The task utilizes the QM9 dataset (Ramakrishnan et al., 2014). Model performance is assessed using the mean absolute error (MAE) between the predicted and ground-truth values.